Cullumber Company has had 4 years of record earnings. Due to this success, the market price of its 500,000 shares of $4 par value common stock has increased from $14 per share to $53. During this period, paid-in capital remained the same at $6,000,000. Retained earnings increased from $4,500,000 to $30,000,000. CEO Don Ames is considering either (1) a 15% stock dividend or (2) a 2-for-1 stock split. He asks you to show the before-and-after effects of each option on (a) retained earnings, (b) total stockholders’ equity, and (c) par value per share.
Answer #1

stock dividend = 500000*.15=75000 shares

Before a 15% stock dividend 2-for-1 stock split
Number of shares after such event 500,000 500,000+ 75000=575,000 500000*2/1=1,000,000
par value $4 $4 $ 4*1/2=$ 2 per share [gets reduced in proportion to stck split
Retained earning 30,000,000

Existing - [Number of shares issued as stock dividend* market value]

30,000,000- [75000*53=$ 3,975,000]

= 26,025,000

No effect [ only shares gets double and par vvalue gets half]

=30,000,000

Additional paid in capital 6,000,000

6,000,000+(75000*(53-4)]

=6,000,000+(75000*49)]

=9675000

6,000,000
Total stockholders equity

[500,000*4]+6,000,000+30,000,000

= 38,000,000

[575000*4]+9675000+26,025,000

=38,000,000

[1,000,000*2]+6,000,000+30,000.000

= 38,000,000